OpenStudy (anonymous):

If f(x) = 2x^2 + 4, find f(3). 22 16 −0.5 40

OpenStudy (anonymous):

put x=3 in the function f(3)=2(3)?^2+4 f(3)=2(9)+4 f(3)=18+4 f(3)=22

OpenStudy (anonymous):

thnks for doing the steps to solving the equation